# Detecting Multimedia Generated by Large AI Models: A Survey

### v1

@@ Abstract
The rapid advancement of Large AI Models (LAIMs), particularly diffusion models and large language models, has marked a new era where AI-generated multimedia is increasingly integrated into various aspects of daily life. Although beneficial in numerous fields, this content presents significant risks, including potential misuse, societal disruptions, and ethical concerns. Consequently, detecting multimedia generated by LAIMs has become crucial, with a marked rise in related research. Despite this, there remains a notable gap in systematic surveys that focus specifically on detecting LAIM-generated multimedia. Addressing this, we provide the first survey to comprehensively cover existing research on detecting multimedia (such as text, images, videos, audio, and multimodal content) created by LAIMs. Specifically, we introduce a novel taxonomy for detection methods, categorized by media modality, and aligned with two perspectives: pure detection (aiming to enhance detection performance) and beyond detection (adding attributes like generalizability, robustness, and interpretability to detectors). Additionally, we have presented a brief overview of generation mechanisms, public datasets, and online detection tools to provide a valuable resource for researchers and practitioners in this field. Furthermore, we identify current challenges in detection and propose directions for future research that address unexplored, ongoing, and emerging issues in detecting multimedia generated by LAIMs. Our aim for this survey is to fill an academic gap and contribute to global AI security efforts, helping to ensure the integrity of information in the digital realm. The project link is https://github.com/Purdue-M2/Detect-LAIM-generated-Multimedia-Survey.
@@ 1.1 Related Works
While there are a few surveys [18, 19, 20, 21, 22] addressing the detection of multimedia generated by LAIMs, like Diffusion Models (DMs) and Large Language Models (LLMs), their scopes are very limited. Specifically, surveys like [18, 19, 20, 21] mainly concentrate on the detection of LAIM-generated text, overlooking other multimedia forms such as images, videos, and audios. Additionally, while these surveys provide insights into detection techniques, they tend to focus solely on detection mechanisms without delving into broader aspects or applications of these technologies beyond detection, which are discussed in our survey.
The survey by Cardenuto et al. [22] falls short of being comprehensive. They present a high-level overview of challenges and opportunities in generation and detection using a broad range of AI models. However, their focus is not specifically on large AI models, leading to an oversight of numerous existing detection works. Additionally, their survey lacks in-depth discussion about datasets in the context of detection. Crucially, the detection methods discussed in [22] are confined to forensic types and do not delve into the specifics of detection models and methodologies, which is one of the central focuses of our survey. They also overlook detection approaches that utilize multimodal data modality, an area our survey encompasses.
In summary, there is currently no comprehensive survey covering the detection of fake media in text, images, videos, audio, and multimodal content, especially from the perspective of large AI models. Our survey attempts to fill this gap by making the following contributions.
@@ 1.2 Contributions
    This is the first comprehensive survey on detecting multimedia generated by LAIMs, covering text, images, videos, audio, and multimodal content. Through this organized and systematic survey, we hope to help push research in this field forward.
@@
    We provide a brief overview of LAIMs, their generation mechanisms, and the content they generate, shedding light on the current status of the cat (generation)-and-mouse (detection) game in this field. Additionally, we present public datasets tailored for detection tasks.
@@
    This survey reviews the most up-to-date detection methodologies for LAIM-generated multimedia. We innovatively categorize them for each modality into two main groups: pure detection and beyond detection. This taxonomy offers a unique perspective that has not been previously explored. Within these two categories, we further classify the methods into more specific subcategories based on their common and distinct characteristics.
@@
    We have conducted thorough examinations of online detection tools. This provides a valuable resource for both researchers and practitioners in the field. Additionally, we pinpoint current challenges faced by recent detection methods and provide directions for future work.
The remained of this survey is organized as follows: In Section 2, we briefly introduce LAIMs, including generation mechanisms and datasets. In Section 3, we classify detection by their functionality and organize them into our newly defined taxonomy. We summarize online detection tools in Section 4. Critical challenges faced by detectors and potential future directions are discussed in Section 5. Finally, we conclude the survey in Section 6.
@@
Modality Dataset Content Link I2O #Real #Generated Source of Real Media Generation Method Year
Text Stu.Essays[23] Essays Link T2T 1,000 6,000 IvyPanda [24] ChatGPT 2023
Writing[23] Essays Link T2T 1,000 6,000 Reddit WritingPrompts [25] ChatGPT 2023
News[23] Essays Link T2T 1,000 6,000 Reuters 50-50 [26] ChatGPT 2023
Paraphrase[27] Essays Link T2T 98,280 163,710 Arxiv, Wikipedia, Theses GPT-3, T5[28] 2022
AA[29] Essays Link T2T 1,064 8,512 News Media GPT-1&2, CTRL[30], GROVER[31] 2020
OUTFOX [32] Essays Link T2T 15,400 15,400 Feedback Prize [33] ChatGPT, GPT-3.5, T5 [28] 2023
MULTITuDE[34] News Link T2T 7,992 66,089 MassiveSumm [35] GPT-3&4, ChatGPT 2023
TuringBench[36] News Link T2T 8,854 159,758 News Media GPT-1&2&3, CTRL [30], GROVER [31] 2021
GPTUnmixed[37] News Link T2T 5,454 5,454 News Media GPT-3.5 2023
GPTMixed[37] News Link T2T 5,032 5,032 News Media GPT-3.5 2023
GPABench[38] Writing Link T2T 150,000 450,000 Arxiv GPT-3.5 2023
HPPT[39] Abstracts Link T2T 6,050 6,050 ACL Anthology [40] ChatGPT 2023
TweepFake [41] Tweets Link T2T 12,786 12,786 GitHub, Twitter GPT-2, RNN, LSTM 2021
SynSciPass [42] Passages Link T2T 99,989 10,485 Scientific papers GPT-2, BLOOM[43] 2022
DFText[44] General Link T2T 154,078 294,381 Reddit, ELI5[45],Yelp, XSum [46] GPT, GLM [47], LLaMA, T5 [28] 2022
HC-Var[48] General Link T2T 90,096 45,000 XSum [46], IMDb, Yelp, FiQA [49] ChatGPT 2023
HC3 [50] General Link T2T 26,903 58,546 FiQA[49], ELI5[45], Meddialog[51] ChatGPT 2023
M4[52] General Link T2T 32,798 89,683 WikiHow [53], Arxiv, Reddit ChatGPT, LLaMA,T5 [28], BLOOM [43] 2023
MixSet[54] General Link T2T 300 3,600 Email[55], BBC News[56], ArXiv
GTP-4, LLaMA2 [57]
2024
InternVid[58] Captions Link V2T 7,000,000 234,000,000 YouTube ViCLIP [58] 2023
Image DFF[59] Face Link T/I2I 30,000 90,000 IMDB-WIKI[60] SDMs, InsightFace [61] 2023
RealFaces [62] Face Link T2I 258 25,800 Prompts SDMs 2023
OHImg [63] Overhead Link T/I2I 6,475 6,675 MapBox [64], Google Maps GLIDE [65], DDPM [66] 2023
Western Blot[67] Biology Link I2I $\sim$14,000 $\sim$24,000 Western Blot DDPM, Pix2pix [68], CycleGAN [69] 2022
Synthbuster[70] General Link T2I - 9,000 Raise-1k[71] DALL·E 2&3, Midjourney, SDMs, GLIDE[65] 2023
GenImage[72] General Link T/I2I 1,331,167 1,350,000 ImageNet SDMs, Midjourney, BigGAN [73], 2023
CIFAKE[74] General Link T2I 60,000 60,000 CIFAR-10 SD-V1.4 2023
AutoSplice[75] General Link T2I 2,273 3,621 Visual News [76] DALL·E 2 2023
DiffusionDB[77] General Link T2I 3,300,000 16,000,000 DiscordChatExporter [78] SD 2023
ArtiFact[79] General Link T/I2I 964,989 1,531,749 COCO, FFHQ [80], LSUN SDMs, DDPM [66],LDM [81], CIPS [82] 2023
HiFi-IFDL[83] General Link T/I2I $\sim$600,000 1,300,000 FFHQ [80], COCO, LSUN DDPM [66], GLIDE [65], LDM [81],GANs 2023
DiffForensics[84] General Link T/I2I 232,000 232,000 LSUN, ImageNet LDM [81], DDPM [66], VQDM [85], ADM [86] 2023
CocoGlide[87] General Link T2I 512 512 COCO GLIDE[65] 2023
LSUNDB[88] General Link T/I2I 250,000 250,000 LSUN DDPM [66], LDM [81], StyleGAN [80] 2023
UniFake[89] General Link T2I 8,000 8,000 LAION-400M [90] LDM [81], GLIDE [65] 2023
REGM [91] General Link T/I2I - 116,000 CelebA [92], LSUN 116 publicly available GMs 2023
DMimage [93] General Link T2I 200,000 200,000 COCO, LSUN LDM [81] 2022
AIGCD[94] General Link T/I2I 360,000 508,500 LSUN, COCO, FFHQ [80] SDMs, GANs,ADM [86],DALL·E 2,GLIDE[65] 2023
DIF [95] General Link T/I2I 84,300 84,300 LAION-5B [96] SDMs, DALL·E 2, GLIDE [65], GANs 2023
Fake2M[97] General Link T/I2I - 2,300,000 CC3M[98] SD-V1.5 [99], IF [100], StyleGAN3 2023
Video DiffHead[101] Face Link I.A2V - 820 CREMA [102] Diffused Heads: build on DDPM 2023
Audio LibriSeVoc[103] Speech Link T2A 13,201 79,206 LibriTTS[104] DiffWave [105], WaveNet [106] 2023
Multi- modal $DGM^{4}$ [14] News Link T/I2T 77,426 152,574 Visual News [76] B-GST [107], StyleCLIP [108], HFGI [109] 2023
COCOFake[110] General Link T/I2T 113,287 566,435 COCO SDMs 2023
@@
TABLE I: Summary of public datasets that are generated by LAIMs. I2O: Input-to-Output, T2T: Text-to-Text, V2T: Video-to-Text, T/I2I: Text-to-Image,Image-to-Image, T2A: Text-to-Audio, I.A2V: (Image conditioned with Audio)-to-Video. Only representative works are listed in ”Source of Real Media” and “Generative Method”.
@@ 2 Generation
$\clubsuit$ Text. Machine-generated text, primarily driven by the advent of LLMs, is increasingly permeating many aspects of our daily lives. The exceptional proficiency of LLMs in understanding, following, and complex reasoning [111] has established their dominance in text creation. Recently, we have witnessed a surge of LLMs including OpenAI GPT-4 [112], Google Gemini [113], Meta LLaMA2 [57] as well as the remarkable performance they have achieved in various tasks, such as News [31, 36], Question Answering [45], Biomedicine [114], Code Generation [115], Tweets [41], and Scientific Writing [42], see Fig. 2 a). More details can be found in [116, 1].
Datasets. The prevalent datasets for LLM-generated text detection are listed in Table I. For example, HC3 [50] stands as one of the initial open-source efforts to compare ChatGPT-generated text with human-written text. Due to its pioneering contributions in this field, the HC3 corpus has been utilized as a valuable resource in numerous subsequent studies. Moreover, Authorship Attribution [29] and TuringBench[36] datasets are proposed to evaluate the attribution capability of detectors. HPPT [39] can benchmark task for detecting ChatGPT-polished texts.
@@
$\clubsuit$ Image. In the challenging task of image synthesis, diffusion models (DMs) [66, 117, 118] have emerged as the new state-of-the-art family of deep generative models. The image generation process in DMs usually contains two processes [66]: a forward process that progressively destroys data by adding noise and a reverse process that learns to generate new data by denoising. More details can be found in [4, 119]. Current research on diffusion models is mostly based on three predominant formulations: denoising diffusion probabilistic models (DDPMs) [66], score-based generative models (SGMs) [117], and stochastic differential equations (Score SDEs) [118]. Building upon them, more advanced models have emerged in image generation, including OpenAI DALL·E 3 [120], Stable Diffusion V2 [81], Google Imagen2 [121], Midjourney [122], Amazon Titan Image Generator [123], and Meta Emu Edit [124], see Fig. 2 b).
Datasets. There are several million-scale generated images detection datasets: GenImage [72], DiffusionDB [77], ArtiFact [79], HiFi-IFDL [83], and Fake2M [97]. DiffusionDB stands out as the first large-scale Image-Prompt pairs dataset, though its image is only generated by Stable Diffusion. GenImage, ArtiFact, HiFi-IFDL, and Fake2M contain images generated by various DMs and GANs. Unique to HiFi-IFDL is its inclusion of high-resolution forgery masks, making it particularly useful for both detection and localization tasks. Refer Table I for more details.

### v4

@@ Abstract
The rapid advancement of Large AI Models (LAIMs), particularly diffusion models and large language models, has marked a new era where AI-generated multimedia is increasingly integrated into various aspects of daily life. Although beneficial in numerous fields, this content presents significant risks, including potential misuse, societal disruptions, and ethical concerns. Consequently, detecting multimedia generated by LAIMs has become crucial, with a marked rise in related research. Despite this, there remains a notable gap in systematic surveys that focus specifically on detecting LAIM-generated multimedia. Addressing this, we provide the first survey to comprehensively cover existing research on detecting multimedia (such as text, images, videos, audio, and multimodal content) created by LAIMs. Specifically, we introduce a novel taxonomy for detection methods, categorized by media modality, and aligned with two perspectives: pure detection (aiming to enhance detection performance) and beyond detection (adding attributes like generalizability, robustness, and interpretability to detectors). Additionally, we have presented a brief overview of generation mechanisms, public datasets, online detection tools, and evaluation metrics to provide a valuable resource for researchers and practitioners in this field. Most importantly, we offer a focused analysis from a social media perspective to highlight their broader societal impact. Furthermore, we identify current challenges in detection and propose directions for future research that address unexplored, ongoing, and emerging issues in detecting multimedia generated by LAIMs. Our aim for this survey is to fill an academic gap and contribute to global AI security efforts, helping to ensure the integrity of information in the digital realm. The project link is https://github.com/Purdue-M2/Detect-LAIM-generated-Multimedia-Survey.
@@ 1.1 Related Works
While there are several surveys [18, 19, 20, 21, 22, 23, 24, 25, 26] addressing the detection of multimedia generated by LAIMs, like Diffusion Models (DMs) and Large Language Models (LLMs), their scopes are very limited. Specifically, surveys like [18, 19, 20, 21, 22] mainly concentrate on the detection of LAIM-generated text, overlooking other multimedia forms such as images, videos, and audio. Additionally, while these surveys provide insights into detection techniques, they tend to focus solely on detection mechanisms without delving into broader aspects or applications of these technologies beyond detection or without discussing real-world challenges like social media adaptation, which are discussed in our survey.
Xu et al. [23] present a focused review on generalizable diffusion-generated image detection. They introduce a dual taxonomy—data-driven and feature-driven—and classify methods into six subcategories based on their design principles. While their survey offers deep insights into generalization and detection robustness for diffusion models, it only addresses a single modality (images) and does not cover other forms of LAIM-generated media such as text, video, or audio.
A few works [24, 25, 26] extend the scope to multimodal detection. Cardenuto et al. [24] offer a high-level overview of generative and detection pipelines but fall short of deep technical analysis. Their scope is not limited to large AI models and omits many recent LAIM-specific detection approaches. Moreover, they do not cover detection datasets in sufficient detail or address multimodal detection settings. Zou et al. [25] and Nguyen-Le et al. [26] both examine multimodal detection, with the former focusing on the transition from Non-MLLM to MLLM-based frameworks and the latter on passive deepfake detection across image, video, audio, and multimodal content. Zou et al. introduce a taxonomy centered on authenticity verification, explainability, and localization, while Nguyen-Le et al. emphasize robustness, generalization, attribution, and real-world deployment. However, both surveys fall short in providing detailed modality-specific analysis, such as concrete challenges, benchmarks, or future directions tailored to each modality.
In summary, current surveys are either not comprehensive in modality coverage or do not address social media–processed multimedia detection in detail, such as datasets, detection methods, challenges, and future directions in a modality-specific manner. Our survey fills this gap by providing the first comprehensive and systematic review of LAIM-generated multimedia detection across five modalities—text, image, video, audio, and multimodal, with an additional dedicated focus on real-world social media settings.
@@ 1.2 Contributions
    This is the first comprehensive survey on detecting multimedia generated by LAIMs, covering text, images, videos, audio, and multimodal content. We review the most up-to-date detection methodologies and propose a novel taxonomy that categorizes methods within each modality into two main groups: pure detection and beyond detection. This taxonomy offers a unique perspective that has not been previously explored. Within these two categories, we further classify the methods into more specific subcategories based on their common and distinct characteristics.
@@
    We provide a brief overview of LAIMs, their generation mechanisms, and the content they generate, shedding light on the current status of the cat (generation)-and-mouse (detection) game in this field. Additionally, we present public datasets and evaluation metrics tailored for detection tasks.
@@
    We are the first to systematically examine the detection of LAIM-generated multimedia in the context of social media, a major source of AI-generated misinformation. Throughout the survey, we review social media-specific datasets across all modalities, analyze detection methods evaluated on in-the-wild content, and identify unique challenges and future directions tailored to social media environments. Given that social media is a primary channel through which AI-generated content spreads and influences public perception, understanding detection in this context is essential for maintaining societal trust.
@@
    We have conducted a comparative summary of online detection tools. This provides a valuable resource for researchers, developers, and stakeholders in the field. Additionally, we pinpoint current challenges faced by recent detection methods and provide directions for future work.
The remainder of this survey is organized as follows: In Section 2, we briefly introduce LAIMs, including generation mechanisms and datasets. In Section 3, we classify detection by their functionality and organize them into our newly defined taxonomy. We summarize online detection tools and evaluation metrics in Section 4. Critical challenges faced by detectors and potential future directions are discussed in Section 5. Finally, we conclude the survey in Section 6.
@@
TABLE I: Summary of public datasets that are generated by LAIMs. I2O: Input-to-Output, T2T: Text-to-Text, V2T: Video-to-Text, T/I2I: Text-to-Image,Image-to-Image, T2A: Text-to-Audio, I.A2V: (Image conditioned with Audio)-to-Video. BM: Benchmark. Note that the "Year" corresponds to the initial public release (e.g., on arXiv), not necessarily the final publication year. Note that not all works in ”Source of Real Media” and “Generative Method” are listed due to space limitations.
Modality Dataset Year BM Content Link I2O #Real #Generated Source of Real Media Generative Method
Text TuringBench[27] 2021 News Link T2T 8,854 159,758 News Media GPT-1&2&3, CTRL [28], GROVER [29]
Paraphrase[30] 2022 Essays Link T2T 98,280 163,710 Arxiv, Wikipedia, Theses GPT-3, T5[31]
SynSciPass [32] 2022 Passages Link T2T 99,989 10,485 Scientific papers GPT-2, BLOOM[33]
MAGE[34] 2023 General Link T2T 154,078 294,381 Reddit, ELI5[35],Yelp, XSum [36] 27 LLMs
Stu.Essays[37] 2023 Essays Link T2T 1,000 6,000 IvyPanda [38] ChatGPT
Writing[37] 2023 Stories Link T2T 1,000 6,000 Reddit WritingPrompts [39] ChatGPT
News[37] 2023 News Link T2T 1,000 6,000 Reuters 50-50 [40] ChatGPT
OUTFOX [41] 2023 Essays Link T2T 15,400 15,400 Feedback Prize [42] ChatGPT, GPT-3.5, T5 [31]
MULTITuDE[43] 2023 News Link T2T 7,992 66,089 MassiveSumm [44] GPT-3&4, ChatGPT
MGTDetect-CoCo[45] 2023 News Link T2T 10,486 10.484 News Outlets GPT-3.5
HPPT[46] 2023 Abstracts Link T2T 6,050 6,050 ACL Anthology [47] ChatGPT
HC-Var[48] 2023 General Link T2T 90,096 45,000 XSum [36], IMDb, Yelp, FiQA [49] ChatGPT
HC3 [50] 2023 General Link T2T 26,903 58,546 FiQA[49], ELI5[35], Meddialog[51] ChatGPT
M4[52] 2023 General Link T2T 32,798 89,683 WikiHow [53], Arxiv, Reddit ChatGPT, LLaMA,T5 [31], BLOOM [33]
F3 [54] 2023 Social Media Link T2T 12,723 27,667 Politifact [55], Snopes [56] GPT3.5
MixSet[57] 2024 General Link T2T 300 3,600 Email[58], BBC News[59], ArXiv
GTP-4, LLaMA2
GPABench[60] 2024 Writing Link T2T 150,000 450,000 Arxiv GPT-3.5
M4GT-Bench [61] 2024 General Link T2T 93,177 110,388 Wikipedia, WikiHow, Reddit, ArXiv, News 10 LLMs
RAID [62] 2024 General Link T2T 14,971 6,287,820 public datasets from 8 domains 11 LLMs
DetectRL [63] 2024 General Link T2T 100,800 134,400
ArXiv, XSum [36],
Writing Prompts [64], Yelp
GPT-3.5, PaLM2 [65], Claude, LLaMA2
MultiSocial [66] 2024 Social Media - T2T  58,000  414,000
Telegram, Twitter,
Gab, Discord, WhatsApp
7 LLMs
SM-D [67] 2024 Social Media - T2T - - Medium, Quora, Reddit Sourced from social media
Image DFF[68] 2023 Face Link T/I2I 30,000 90,000 IMDB-WIKI[69] SDMs, InsightFace [70]
RealFaces [71] 2023 Face Link T2I 258 25,800 Prompts SDMs
OHImg [72] 2023 Overhead Link T/I2I 6,475 6,675 MapBox [73], Google Maps GLIDE [74], DDPM [75]
Western Blot[76] 2022 Biology Link I2I $\sim$14,000 $\sim$24,000 Western Blot DDPM, Pix2pix [77], CycleGAN [78]
Synthbuster[79] 2023 General Link T2I - 9,000 Raise-1k[80] DALL·E 2&3, Midjourney, SDMs, GLIDE[74]
GenImage[81] 2023 General Link T/I2I 1,331,167 1,350,000 ImageNet SDMs, Midjourney, BigGAN [82],
CIFAKE[83] 2023 General Link T2I 60,000 60,000 CIFAR-10 SD-V1.4
AutoSplice[84] 2023 General Link T2I 2,273 3,621 Visual News [85] DALL·E 2
DiffusionDB[86] 2023 General Link T2I 3,300,000 16,000,000 DiscordChatExporter [87] SD
ArtiFact[88] 2023 General Link T/I2I 964,989 1,531,749 COCO, FFHQ [89], LSUN SDMs, DDPM [75],LDM [90], CIPS [91]
HiFi-IFDL[92] 2023 General Link T/I2I $\sim$600,000 1,300,000 FFHQ [89], COCO, LSUN DDPM [75], GLIDE [74], LDM [90],GANs
DiffForensics[93] 2023 General Link T/I2I 232,000 232,000 LSUN, ImageNet LDM [90], DDPM [75], VQDM [94], ADM [95]
CocoGlide[96] 2023 General Link T2I 512 512 COCO GLIDE[74]
LSUNDB[97] 2023 General Link T/I2I 250,000 250,000 LSUN DDPM [75], LDM [90], StyleGAN [89]
UniFake[98] 2023 General Link T2I 8,000 8,000 LAION-400M [99] LDM [90], GLIDE [74]
REGM [100] 2023 General Link T/I2I - 116,000 CelebA [101], LSUN 116 publicly available GMs
DMimage [102] 2022 General Link T2I 200,000 200,000 COCO, LSUN LDM [90]
AIGCD[103] 2023 General Link T/I2I 360,000 508,500 LSUN, COCO, FFHQ [89] SDMs, GANs,ADM [95],DALL·E 2,GLIDE[74]
DIF [104] 2023 General Link T/I2I 84,300 84,300 LAION-5B [105] SDMs, DALL·E 2, GLIDE [74], GANs
Fake2M[106] 2023 General Link T/I2I - 2,300,000 CC3M[107] SD-V1.5 [108], IF [109], StyleGAN3
SID-Set [110] 2024 Social Media Link T/I2I 100,000 200,000 COCO, Flickr30k, MagicBrush FLUX [111]
Chameleon [112] 2024 Social Media Link T/I2I 14,863 11,170 Unsplash [113] GANs, SDMs, DALL·E 2,GLIDE[74]
DF40 [114] 2024 Face Link T/I2I - $\sim$ 1,000,000 FF++, CDF, FFHQ, CelebA SDMS, GANs, Midjourney, DDPM
FakeBench [115] 2024 General Link T/I2I 3,000 3,000 10 Public Datasets 10 Generative Models
AI-Face [116] 2024 Face Link T/I2I 400,885 1,245,660 6 Public datasets SDMs, GANs, Midjourney, IF
Video WildDeepfake[117] 2021 Face Link - 3,805 3,509 Social Media Social Media
DiffHead[118] 2023 Face Link I.A2V - 820 CREMA [119] Diffused Heads: build on DDPM
DVF[120] 2024 General Link I/T2V 2,750 3,938 Internvid [121], Youtube-8M [122] 8 Diffusion Models
GenVideo[123] 2024 General Link I/T2V 1,223,511 1,078,838 Kinetics-400 [124], Youku-mPLUG [125], MSR-VTT [126] 20 Generative Models
GenVidBench[127] 2025 General Link I/T2V 33,931 110,400 Vript[128], HD-VG-130M [129] 8 Generative Models
PDID[130] 2024 Social Media Link - - - Social Media Social Media
Audio In-the-Wild[131] 2022 Speech Link - 20.7 hours 17.2 hours Social Media, Video Streaming Platforms Social Media, Video Streaming Platforms
LibriSeVoc[132] 2023 Speech Link T2A 13,201 79,206 LibriTTS[133] DiffWave [134], WaveNet [135]
SONAR[136] 2024 Speech Link T2A - 2,274 LibriTTS[133] OpenAI [137], Seed-TTS [138], AudioGen [139]
ASVspoof 2024[140] 2024 Speech Link T/A2A $\sim$289,527 $\sim$1,211,186 MLS-English [141] 32 Manipulation Methods
Multi- modal $DGM^{4}$ [14] 2023 News Link T/I2T 77,426 152,574 Visual News [85] B-GST [142], StyleCLIP [143], HFGI [144]
COCOFake[145] 2023 General Link T/I2T 113,287 566,435 COCO SDMs
AV-Deepfake1M[146] 2023 Face Link
T2A
I2I
286,721 860,039 Voxceleb2 [147], VITS [148], YourTTS [149], TalkLip [150]
$D^{3}$ [151] 2024 General Link T2I $\sim$2,300,000 $\sim$9,200,000 LAION-400M SDMs, IF
$M^{3}A$[152] 2024 - News -
T2T/I/V/A
T/I/V/A2T
708,425 6,566,386 60 News Outlets LLaMA2, GPT-4, GLIDE [74], SD, Tango [153]
LOKI[154] 2024 General Link
T2T/I/V/A
T/I/V/A2T
$\sim$9,000 $\sim$9,000 21 Public Datasets 43 Generative Models
MMFakeBench[155] 2024 Social Media Link
T2T/I
I2I
- $\sim$11,000
MS-COCO, VisualNews,
Reddit,FEVER [156]
GPT-3.5, SD-XL, DALL·E 3, Midjourney
Deepfake-Eval[157] 2024 Social Media Link
T2I/A/V
I2I/V
3,390 2,441 Social Media Social Media
ILLUSION[158] 2025 General Link
T2A/I
I2I
139,740 1,232,246
CelebV-Text [159], COCO,
MusicCaps [160], Social Media
28 Generative Methods
@@ 2 Generation
$\clubsuit$ Text. Machine-generated text, primarily driven by the advent of LLMs, is increasingly permeating many aspects of our daily lives. The exceptional proficiency of LLMs in understanding, following, and complex reasoning [161] has established their dominance in text creation. Recently, we have witnessed a surge of LLMs including OpenAI GPT-4 [Achiam2023GPT4TR], Google Gemini [162], Meta LLaMA2 [163] as well as the remarkable performance they have achieved in various tasks, such as News [29, 27], Question Answering [35], Biomedicine [164], Code Generation [165], Tweets [166], and Scientific Writing [32], see Fig. 2 a). More details can be found in [167, 1].
LAIM Generated Datasets (Recommended). The datasets for LLM-generated text detection are listed in Table I. Recent efforts have introduced large-scale and challenging benchmarks to evaluate detector robustness under realistic and adversarial settings. For instance, RAID [62] is the largest and most comprehensive benchmark to date, containing over 6 million texts generated by 11 LLMs across 8 domains, 4 decoding strategies, and 11 types of adversarial attacks. It supports fine-grained evaluation of generalization and robustness by simulating real-world perturbations such as paraphrasing, machine translation, and sampling noise. DetectRL [63] focuses on real-world detection by integrating reinforcement learning strategies and human-in-the-loop edits to simulate text that mimics post-edited or polished LLM outputs. MAGE [34] presents a diverse benchmark combining 27 LLMs and 7 writing tasks, including both creative and factual domains, enabling a broad analysis of generalization performance across model families and genres.
Social Media. F3 [54] introduces a novel dataset to study both generation and detection of disinformation across news articles and social posts, leveraging perturbation- and paraphrase-based generation with models like GPT-3.5. Similarly, MultiSocial [66] fills a critical gap by offering the first multilingual, multi-platform benchmark of social media texts, covering 22 languages and 5 platforms. Its focus on informal, short-form user-generated content makes it particularly valuable for evaluating cross-lingual and domain-specific detection robustness.
Non-LAIM Generated Datasets. AA [168], TweepFake [166], and GPT-2 Output [169] are widely used datasets produced by earlier language models (e.g., n-gram models, RNNs, LSTMs, or smaller transformer models like GPT-2). These datasets typically include machine-generated text that exhibits limited fluency, shallow coherence, and frequent repetition, especially in longer sequences. OpenAI’s analysis of GPT-2 outputs revealed notable distributional artifacts, such as the underuse of proper nouns and overuse of pronouns and generic expressions, which early detectors could effectively exploit [169]. In contrast, text generated by LLMs, as seen in datasets like RAID [62] or DetectRL [63], is significantly more fluent, coherent, and contextually accurate.
@@
$\clubsuit$ Image. In the challenging task of image synthesis, diffusion models (DMs) [75, 170, 171] have emerged as the new state-of-the-art family of deep generative models. The image generation process in DMs usually contains two processes [75]: a forward process that progressively destroys data by adding noise and a reverse process that learns to generate new data by denoising. More details can be found in [4, 172]. Current research on diffusion models is mostly based on three predominant formulations: denoising diffusion probabilistic models (DDPMs) [75], score-based generative models (SGMs) [170], and stochastic differential equations (Score SDEs) [171]. Building upon them, more advanced models have emerged in image generation, including OpenAI DALL·E 3 [173], Stable Diffusion 3.5 [174], Google Imagen 3 [175], Midjourney 7 [176], Amazon Titan Image Generator 2 [177], and Meta Emu Edit [178], see Fig. 2 b).
LAIM Generated Datasets (Recommended). There are several million-scale LAIM-generated image detection datasets, see Table I. Among them, GenImage[81] offers over 2.6M real and fake images across 1000 ImageNet categories, generated by both GANs and diffusion models. It emphasizes diverse content and proposes tasks such as cross-generator and degraded-image detection to simulate real-world challenges. DiffusionDB[86], in contrast, uniquely pairs 14M images with 1.8M user-written prompts, revealing generation patterns and biases in human-guided diffusion-based generation, though it is limited to Stable Diffusion outputs. AI-Face [116] introduces the first million-scale demographically annotated dataset of AI-generated faces, incorporating 37 generation methods (deepfakes, GANs, and DMs) and annotated gender, age, and skin tone. It also provides a comprehensive fairness benchmark for evaluating detection disparities across demographic groups. DF40[114] pushes realism and diversity further by including 40 modern face synthesis techniques spanning face-swapping, reenactment, editing, and full-synthesis, reflecting the heterogeneity of real-world manipulations and bridging the gap between academic and practical detection scenarios.
Social Media. SID-Set [110] is a comprehensive dataset with 300K real, synthetic, and tampered images annotated for detection, localization, and explanation. It emphasizes scene complexity and visual plausibility, targeting real-world misinformation scenarios. Chameleon [112] complements this by curating images that consistently fool human annotators, representing high-resolution, semantically rich content from categories such as humans, animals, and objects. Detection models struggle significantly on this benchmark, highlighting current limitations.
Non-LAIM Generated Datasets. FaceForensics++ (FF++) [179], DeepfakeDetection (DFD) [180], Deepfake Detection Challenge (DFDC) [181], and Celeb-DF [182] are four most representative deepfake video datasets. Although originally constructed as video datasets, they are often used for frame-level detection; therefore, we also categorize them under image-based datasets. They are created by early manipulation methods, such as graphics-based [183, 184], autoencoder-based [185], and GAN-based [186]. These methods typically rely on paired source-target videos, facial tracking, and region blending to produce partially manipulated videos that often leave visible artifacts, such as blending boundaries around the face, color mismatches between the manipulated region and surrounding skin, temporal flickering across frames, and unnatural mouth or eye movements. In contrast, diffusion models generate content entirely from scratch, using text or image prompts. This enables highly customized and diverse outputs in terms of style, identity, and context. Unlike traditional deepfake methods, diffusion models do not require explicit source-target video pairs and often produce artifact-free content, making detection significantly more challenging.